# What Happens During the Loss Plateau? Understanding Abrupt Learning in Transformers

### v1

@@ Appendix C Results for REPEAT2 and REPEAT4
[…]
The training dynamics for REPEAT2 are given in Figure 19(a). We find that similar to REPEAT1, the training loss does not exhibit any plateau. Moreover, the repetition frequency $\rho$ and metric $\alpha_{1}$ increase rapidly to $\approx 1.0$ early on in training. We measure the average cosine similarity for hidden states for repetitive blocks of the output sequence (see Figure 19(b)).
@@
[…]
Similar results are observed for REPEAT4 as well; for this case we measure the average cosine similarity for hidden states for repetitive blocks of the output sequence (see Figure 20(b)).
@@
Appendix D Varying Configurations
We demonstrate below that abrupt learning, representation collapse and repetition bias occur across model hyperparameter variations for the MWS task (Figures 21, 22, 24, 23 and 25). We also show that training using SGD instead of Adam also demonstrates similar abrupt learning characteristics (Figure 26).
Figure 21: Number of Layers $(L)$. We show that abrupt learning, representation collapse in the last layer and repetitions occur for multi (2, 4, 6)–layer models as well.
@@ Appendix D Varying Configurations
Figure 23: Number of attention heads $(H)$. We show that abrupt learning with representation collapse and repetition bias occurs in 1-layer multi-attention head models as well.
Figure 24: Embedding dimension $(d)$. We show that abrupt learning with representation collapse and repetition bias occurs in 1-layer 1-head models with different embedding dimension. Note that the convergence is delayed for models with smaller values of $d=64,128.$
@@
Figure 26: SGD instead of Adam for loss optimization. We show that abrupt learning is not limited to Adam optimizer, and occurs with SGD $(\eta=0.1)$ as well. We chose this value of $\eta$ since smaller values typically lead to much longer periods of little decrease in loss, without increase in accuracy.
Appendix E Additional Figures
@@
Figure 30: Representation collapse at different Pythia pretraining checkpoints; inference with random sampling.

### v2

@@ Appendix D Results for REPEAT2, REPEAT4
$y_{i}=\begin{cases}x_{1}&1\leq i\leq 8\\ (x_{1}+1)\,{\rm mod}\,\,17&9\leq i\leq 16\end{cases}$
The training dynamics for REPEAT2 are given in Figure˜21(a). We find that similar to REPEAT1, the training loss does not exhibit any plateau. Moreover, the repetition frequency $\rho$ and metric $\alpha_{1}$ increase rapidly to $\approx 1.0$ early on in training. We measure the average cosine similarity for hidden states for repetitive blocks of the output sequence (see Figure˜21(b)).
@@
$y_{i}=\begin{cases}x_{1}&1\leq i\leq 4\\ (x_{1}+1)\,{\rm mod}\,\,17&5\leq i\leq 8\\ (x_{1}+2)\,{\rm mod}\,\,17&9\leq i\leq 12\\ (x_{1}+3)\,{\rm mod}\,\,17&13\leq i\leq 16\end{cases}$
Similar results are observed for REPEAT4 as well (Figure˜22(a)); for this case we measure the average cosine similarity for hidden states for repetitive blocks of the output sequence (see Figure˜22(b)).
@@
Appendix E Additional LLM Results
Figure 23: Representation Collapse, Repetitions during Pythia Pretraining (GSM-8K [13]). Representation collapse at different Pythia pretraining checkpoints evaluated on the GSM-8K test dataset; tokens generated via greedy decoding.
Figure 24: Representation Collapse, Repetitions during Pythia Pretraining (ARC-Challenge [12]). Representation collapse at different Pythia pretraining checkpoints evaluated on the ARC-Challenge test dataset; tokens generated via greedy decoding.
Figure 25: Representation collapse during Pythia pretraining (inference with random sampling).
Appendix F Varying Model Configurations
Figure 26: Number of Layers $(L)$. Abrupt learning, representation collapse in the last layer and repetitions for multi–layer models.
Figure 27: Number of attention heads $(H)$. Representation collapse and repetition bias in 1-layer multi-attention head models (embedding dimension is fixed at $d=256$ for all cases).
Figure 28: Embedding dimension $(d)$. Abrupt learning with representation collapse and repetition bias in 1-layer 1-head models with different embedding dimensions.
@@ Appendix F Varying Model Configurations
Appendix G Varying Optimization Setups
Figure 31: Using SGD for training. We show that abrupt learning is not limited to Adam optimizer, and occurs with SGD $(\eta=0.1)$ as well. We chose this value of $\eta$ since smaller values typically lead to much longer periods of little decrease in loss, without increase in accuracy.
Figure 32: Using Muon for training, with Adam for embeddings, LM head and bias parameters. We use learning rate $0.02$ for Muon, and 11e–44 for Adam. Note that while the plateau duration decreases significantly compared to using only Adam (Figure˜2(a)), a plateau of duration $\sim 40$ steps still occurs.
Figure 33: Using Muon for training, with AdamW for embeddings, LM head and bias parameters. We use learning rate $0.02$ for Muon, and 11e–44 for AdamW (weight decay=0.01=0.01).
Figure 34: Training Dynamics with Learning Rate Warmup. The learning rate increases linearly from 0 to 11e–44 during training steps $[0,t]$.
Figure 35: Training Dynamics with Learning Rate Warmup + Cosine Decay. The learning rate increases linearly from 0 to 11e–44 during training steps $[0,t]$, and decreases as a cosine function during steps $[t+1,400].$
Figure 36: Training Dynamics with different Initialization Scales. We find that smaller initialization scale $(\alpha<1)$ leads to a longer plateau, and cosine similarity being very close to $1.0$ throughout the plateau. With $\alpha=10$, the representation collapse is not as strong, reaching a peak value of $\approx 0.6$. We smoothen the repetition frequency curve with window of size 2020 for presentation clarity, using np.convolve(repeat_freq, np.ones(20)/20, mode=’same’).
Figure 37: Training Dynamics with different Learning Rates.
Figure 38: Training Dynamics with varying Batch Size.
Figure 39: Training Dynamics with AdamW, varying weight decay parameter $(\lambda)$.
Figure 40: Training Dynamics with $\ell_{2}$ regularization, varying regularization parameter $(\lambda)$.
Appendix H Additional Figures
@@
Appendix I Additional Related Work
Techniques from statistical physics have also been used towards understanding initial loss plateaus in neural net training [43, 24]; they work in a 2-layer teacher-student setup, where the second layer is fixed during training, and use order parameters to study training. They show that there is a permutation symmetry in the weight vectors of the first layer during the early plateau stage, and exiting this state leads to drop in loss. Singular learning theory [22] has also been used to explain stagewise learning dynamics in Transformers; they estimate the ‘Local Learning Coefficient’ (LLC) during training to quantify degeneracy in the loss landscape, and consequently explain the learning dynamics. The interplay of simplicity bias and Transformer learning dynamics has also been studied recently in [42, 5]. [42] show that Transformers progressively learn higher-order (‘many-body’) interactions among tokens in the sequence. In [5] authors show that neural nets learn to use lower-order moments of data early in training via interventions on test data, and show an equivalence between $n$-gram statistics and embedding moments for discrete domains.